### v1

@@ 2. Generate a sample initial state $x_{0}^{(k)}$ with $\lVert x_{0}^{(k)}\rVert\leq L_{0}$;
Based on the results of Theorem 1 and Theorem 31, we conclude that, with appropriate choices of $l$, $r$, and $n$, the estimates $\hat{\Sigma}_{K}$ and ^$\hat{\nabla}C(K)$ from Algorithm 1 can achieve any desired accuracy with a corresponding desired probability. These estimates then replace their model-based counterparts in model-free policy gradient methods. The error bounds established in this section serve as a foundation for the convergence guarantees of both model-free gradient descent and natural policy gradient methods, which are discussed in the next sections.
@@ 4.2 Model-free Gradient Estimation with Variance Reduction
In the previous section, we employed zeroth-order optimization to estimate the gradient of the cost at the current value of the policy. However, this approach often suffers from high variance, resulting in a slow learning process [12]. Using a baseline is a common variance reduction technique in policy gradient methods [9, 21]. In this section, we propose employing the finite-horizon cost function as a baseline and show its performance improvement. For a state-dependent baseline function $b(x)$, the estimated gradient with variance reduction ^$\hat{\nabla}C_{v}(K)$ is expressed as
(34) ^$\hat{\nabla}C_{v}(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}(\hat{C}_{k}-b(x_{0}^{(k)}))U_{k},$
where $U_{k},x_{0}^{(k)},\hat{C}_{k},r$ are the same as those defined in Algorithm 1. The estimated gradient ^$\hat{\nabla}C_{v}(K)$ satisfies:
(35) $\begin{split}\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C_{v}(K)]=\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C(K)]-\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}b(x_{0}^{(k)})U_{k}].\end{split}$
@@
Next, we determine the optimal function $b^{*}$, which minimizes the variance of the sample $Z_{k}^{V}:=(\hat{C}_{k}-b(x_{0}^{(k)}))U_{k}$. The norm of the variance of $Z_{k}^{V}$ is given by:
(36) $\lVert Var[Z_{k}^{V}]\rVert_{F}=\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[\left\lVert(\hat{C}_{k}-b(x_{0}^{(k)}))U_{k}\right\rVert_{F}^{2}\right]-\left\lVert\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[(\hat{C}_{k}-b(x_{0}^{(k)}))U_{k}\right]\right\rVert_{F}^{2}.$
@@
Consequently, the second term in (36) becomes independent of the choice of $b$, leaving only the first term in (36) to be minimized. To minimize the first term, the baseline function should be chosen to reduce the mean square error of $\hat{C}_{k}$. Thus, the optimal baseline $b^{*}$ is given by:
(37) $b^{*}(x_{0})=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{w_{t},U_{k}}\left[x_{t}^{~{}\top}(Q+(\hat{K}_{i}+U_{k})^{\top}R(\hat{K}_{i}+U_{k}))x_{t}\right].$
Since $U_{k}$, which was also already defined in Algorithm 1, enters nonlinearly in (37), it is hard to compute the expectation and thus to obtain a closed-form expression for $b^{*}(x_{0})$. Instead, a suboptimal solution $b_{s}(x_{0})$ is used:
(38) $b_{s}(x_{0})=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{w_{t}}\left[x_{t}^{~{}\top}(Q+\hat{K}^{\top}_{i}R\hat{K}_{i})x_{t}\right].$
While the proposed approximation is simply driven by the need to simplify the computation of the expected value in (37), we will show later (Theorem 3) that using $b_{s}$ yields a provable reduction in the required number of rollouts compared to the standard gradient computation from the previous section.
@@
Algorithm 2 Baseline function estimation $b_{s}(x_{0})$ for variance reduction of ^$\hat{\nabla}C(K)$
@@
      1. Excite the closed-loop system with $u_{t}=\hat{K}_{i}x_{t}$ for $l$-steps starting from $x_{0}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$;
      2. Collect the empirical finite-horizon cost estimate $\hat{C}^{V}_{k}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~{}\top}(Q+\hat{K}^{\top}_{i}R\hat{K}_{i})x_{t}^{(k)}$;
@@ 2. Estimate the baseline function $\hat{b}_{s}(x_{0}^{(k)})$ using Algorithm 2
      3. Generate a sample gain matrix $\hat{K}_{k}=K+U_{k}$, where $U_{k}$ is drawn uniformly at random over matrices of compatible dimensions with Frobenius norm $r$;
      4. Excite the closed-loop system with $u_{t}^{(k)}=\hat{K}_{k}x_{t}^{(k)}$ for $l$-steps starting from $x_{0}^{(k)}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$;
      5. Collect the empirical finite-horizon cost estimate $\hat{C}_{k}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~{}\top}(Q+\hat{K}^{\top}_{k}R\hat{K}_{k})x_{t}^{(k)}$
@@
  return  Gradient estimate ^$\hat{\nabla}C(K):=\frac{1}{n_{b}}\sum^{n_{b}}_{k=1}\frac{n_{x}n_{u}}{r^{2}}(\hat{C}_{k}-\hat{b}_{s}(x_{0}^{(k)}))U_{k}$
Compared with Algorithm 1, Algorithm 3 introduces an additional step involving the baseline function to reduce the variance of the estimated gradient. The following theorem shows that this modification of the algorithm brings a provable improvement in the sample complexity required to estimate the gradient with a given accuracy: {Theorem}[Performance Improvement] Given the same tolerance $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$ and probability $\delta=(1-\delta_{d})(1-\delta_{n})\delta_{x}$ introduced in Theorem 1, for a given $K\in\mathcal{S}$, then the estimated gradient ^$\hat{\nabla}C(K)$ from Algorithm 3 enjoys the following bound:
(39) {^$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq\delta,$
@@
(40a) $r$ $\leq r_{\max}(C(K),\epsilon_{r}),l\geq l_{\min}(C(K),\epsilon_{l}),$
(40b) $n_{b}$ $\geq\max(N_{1},N_{3}(\hat{b}_{s}(x_{0}^{(k)}))$
where the detailed expression of $N_{3}$ is given in (105) in Appendix C.7 and $r_{\max}$, $l_{\min}$ and $N_{1}$ were introduced in Theorem 1.
Moreover, let the desired probability $\delta_{v}\in(0,1)$ be decomposed as $\delta_{v}=(1-\tilde{\delta}_{v})\tilde{\delta}_{x}$. If the probability $\delta_{x}$ satisfies $\delta_{x}\geq\underline{\delta}_{x}(\Sigma_{w})$, and the number of rollouts $n_{v}$ to estimate the baseline function (as defined in Algorithm 2) satisfies
(41) $n_{v}\geq\tilde{n}_{\min}(C(K),l,\tilde{\delta}_{v},\tilde{\delta}_{x}),$
@@
(42) $\mathbb{P}\left\{N_{2}\geq N_{3}\right\}\geq\delta_{v},$
@@
where $N_{2}$ was first introduced in (31b) in Theorem 1 and the detailed expressions of $\underline{\delta}_{x},\tilde{n}_{\min}$ are provided in (103) and (110) in Appendix C.7. The proof of Theorem 3 is given in Appendix C.7. The probability decompositions in Theorem 3 include:
  • $\tilde{\delta}_{x}$: the probability that the sample remains bounded.
  • $\tilde{\delta}_{v}$: the probability associated with the concentration of the sample.
@@
Based on Theorem 3, we can compare the performance of Algorithm 1 without variance reduction and Algorithm 3 with variance reduction. The exploration radius $r_{\max}$ and rollout length $l_{\min}$ (40a) remain the same for both algorithms. The number of rollouts (40b) is determined by the maximum of two arguments. $N_{1}$ remains unchanged across the two algorithms, as it depends on the smoothing function (see (66))and the original function. From the second part of Theorem 3 ((42)), it then follows that the required number of rollouts is at least non-increasing when the baseline function estimates are sufficiently accurate. As noted in Remark 31, $\delta_{x}$ represents the probability that the data samples remain bounded. Increasing $\delta_{x}$ leads to a larger bound for the data samples. This upper bound is utilized in the matrix concentration inequality to determine the required number of rollouts. From (93a) and (93b), we observe that, for large $\Sigma_{w}$ and thus large noise, the required number of rollouts is determined by $N_{2}$ (in Algorithm 1) and $N_{3}$ (in Algorithm 3). In other words, the proposed variance reduction gives a provable reduction of the number of rollouts (and thus of samples) for scenarios with large noise, which is intuitive and expected.
4.3 Gradient Descent with Adaptive Step Size
@@
Building on the gradient uncertainty quantified in the previous section, we now proceed to study the convergence of the model-free gradient descent algorithm. {Theorem}[Model-free Gradient Descent with Adaptive Step Size] Suppose the initial $\hat{K}_{0}\in\mathcal{S}$, and consider the gradient descent with adaptive step size:
(43) $\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i}),\quad\forall i\in\mathbb{Z}_{+},$
@@
where ^$\hat{\nabla}C(\hat{K}_{i})$ is the gradient estimate from Algorithm 1 and $0<\eta_{i}\leq h_{\mathrm{GD}}(C(\hat{K}_{i}))$, where the function $h_{\mathrm{GD}}$ was introduced in Theorem 3.1. Given any accuracy $\epsilon>0$, probability $\delta\in(0,1)$ and $\sigma\in(0,1)$, if the estimation error of the gradient ^$\hat{\nabla}C(\hat{K}_{i})$ satisfies:
@@
{^$\mathbb{P}\left\{\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{h_{C}(C(\hat{K}_{i}))\lVert\Sigma_{K^{*}}\rVert}\right\}\geq\delta,$
@@
where $h_{C}$ was introduced in Lemma 2.2, then for any $C(\hat{K}_{i})\geq C(K^{*})+\epsilon$, the following inequality holds:
$\mathbb{P}\left\{C(\hat{K}_{i+1})-C(K^{*})\leq\gamma_{\mathrm{GD}}(C(\hat{K}_{i})-C(K^{*}))\Big{|}C(\hat{K}_{i})\right\}\geq\delta,$
@@
As a result, the gradient descent method enjoys the following performance bound:
$\mathbb{P}\left\{\min_{i\in[0,n_{\mathrm{GD}}]}C(\hat{K}_{i})-C(K^{*})\leq\epsilon\right\}\geq\delta^{n_{\mathrm{GD}}},$
if the number of iterations $n_{\mathrm{GD}}$ satisfies:
$n_{\mathrm{GD}}\geq\frac{\lVert\Sigma_{K^{*}}\rVert}{2(1-\sigma)\eta_{\mathrm{GD}}\lambda_{1}(R)\lambda^{2}_{1}(\Sigma_{w})}\log\frac{C(\hat{K}_{0})-C(K^{*})}{\epsilon},$
where ${\eta_{\mathrm{GD}}}:=\min\limits_{i\in[0,n_{\mathrm{GD}}]}\eta_{i}$.
Algorithm 4 Model-free gradient descent with adaptive step size

### v2

@@ 2. Generate a sample initial state $x_{0}^{(k)}$ with $\lVert x_{0}^{(k)}\rVert\leq L_{0}$;
Based on the results of Theorem 4 and Theorem 5, we conclude that, with appropriate choices of $l$, $r$, and $n$, the estimates $\hat{\Sigma}_{K}$ and $\hat{\nabla}C(K)$ from Algorithm 1 can achieve any desired accuracy with a corresponding desired probability. These estimates then replace their model-based counterparts in model-free policy gradient methods. The error bounds established in this section serve as a foundation for the convergence guarantees of both model-free policy gradient descent and natural policy gradient methods, which are discussed in the next sections.
@@ 4.2 Model-free Gradient Estimation with Variance Reduction
In the previous section, we employed zeroth-order optimization to estimate the gradient of the cost at the current value of the policy. However, this approach often suffers from high variance, resulting in a slow learning process [48]. Using a baseline is a common variance reduction technique in policy gradient methods [49, 50]. In this section, we propose employing the finite-horizon cost function as a baseline and show its performance improvement. For a state-dependent baseline function $b(x)$, the estimated gradient with variance reduction $\hat{\nabla}C_{v}(K)$ is expressed as
$\hat{\nabla}C_{v}(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}(\hat{C}_{\bar{K}_{k}}-b(x_{0}^{(k)}))U_{k},$ (33)
where $U_{k},x_{0}^{(k)},\hat{C}_{\bar{K}_{k}},r$ are the same as those defined in Algorithm 1. The estimated gradient $\hat{\nabla}C_{v}(K)$ satisfies:
$\begin{split}\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C_{v}(K)]=\mathbb{E}_{x_{0}^{(k)},U_{k}}[\hat{\nabla}C(K)]-\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\mathbb{E}_{x_{0}^{(k)},U_{k}}[b(x_{0}^{(k)})U_{k}].\end{split}$ (34)
@@
The trace of the covariance matrix of vectorized matrix $vec(Z_{k}^{V})$ is given as:
$\mathrm{Tr}(\mathrm{Var}(vec(Z_{k}^{V})))=\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[\left\lVert Z_{k}^{V}\right\rVert_{F}^{2}\right]-\left\lVert\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[Z_{k}^{V}\right]\right\rVert_{F}^{2}.$ (35)
Next, we determine the optimal function $b^{*}$, which minimizes the trace of the variance matrix above:
$\mathrm{Tr}(\mathrm{Var}(vec(Z_{k}^{V})))=\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[\left\lVert(\hat{C}_{\bar{K}_{k}}-b(x_{0}^{(k)}))U_{k}\right\rVert_{F}^{2}\right]-\left\lVert\mathbb{E}_{x_{0}^{(k)},U_{k}}\left[(\hat{C}_{\bar{K}_{k}}-b(x_{0}^{(k)}))U_{k}\right]\right\rVert_{F}^{2}.$ (36)
@@
Consequently, the second term in (36) becomes independent of the choice of $b$, leaving only the first term in (36) to be minimized. To minimize the first term, the baseline function should be chosen to reduce the mean square error of $\hat{C}_{\bar{K}_{k}}$. Thus, the optimal baseline $b^{*}$ is given by:
$b^{*}(x_{0})=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{w_{t},U_{k}}\left[x_{t}^{~\top}(Q+({K}+U_{k})^{\top}R({K}+U_{k}))x_{t}\right].$ (37)
Since $U_{k}$, which was also already defined in Algorithm 1, enters nonlinearly in (37), it is hard to compute the expectation and thus to obtain a closed-form expression for $b^{*}(x_{0})$. While it is possible to estimate it directly from data, the lack of a closed-form solution makes it challenging to establish explicit performance improvement guarantees. Instead, a suboptimal solution $b_{s}$ is used:
$b_{s}(x_{0})=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{w_{t}}\left[x_{t}^{~\top}(Q+{K}^{\top}R{K})x_{t}\right],$ (38)
The rationale for this choice is to simplify the computation of the expected value in (37) so that the performance improvement can be provided. In the next part, we will demonstrate (Theorem 6) that using $b_{s}$ leads to a provable reduction in the number of rollouts required, compared to the standard gradient computation discussed in the previous section.
@@
Algorithm 2 Baseline function estimation $b_{s}(x_{0})$ for variance reduction
@@
   1. Excite the closed-loop system with $u_{t}=Kx_{t}$ for $l$-steps starting from $x_{0}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$;
   2. Collect the empirical finite-horizon cost estimate $\hat{C}^{V}_{k}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+{K}^{\top}R{K})x_{t}^{(k)}$;
@@ 2. Estimate the baseline function $\hat{b}_{s}(x_{0}^{(k)})$ using Algorithm 2
   3. Generate a sample gain matrix $\bar{K}_{k}=K+U_{k}$, where $U_{k}$ is drawn uniformly at random over matrices of compatible dimensions with Frobenius norm $r$;
   4. Excite the closed-loop system with $u_{t}^{(k)}=\bar{K}_{k}x_{t}^{(k)}$ for $l$-steps starting from $x_{0}^{(k)}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$;
   5. Collect the empirical finite-horizon cost estimate $\hat{C}_{\bar{K}_{k}}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+\bar{K}^{\top}_{k}R\bar{K}_{k})x_{t}^{(k)}$
@@
return  Gradient estimate $\hat{\nabla}C(K):=\frac{1}{n_{b}}\sum^{n_{b}}_{k=1}\frac{n_{x}n_{u}}{r^{2}}(\hat{C}_{\bar{K}_{k}}-\hat{b}_{s}(x_{0}^{(k)}))U_{k}$
Compared with Algorithm 1, Algorithm 3 introduces an additional step involving the baseline function to reduce the variance of the estimated gradient. The following theorem establishes that this modification leads to a provable improvement in the sample complexity required to achieve a given gradient accuracy with high probability. Here, the overall probability is expressed as two parts: $\tilde{\delta}_{x}$, representing the probability that the samples remain bounded, and $\tilde{\delta}_{v}$, capturing the concentration of the sample.
Theorem 6 (Performance Improvement)
Given the same tolerance $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$ and probability $\delta=1-(1-\delta_{d})(1-\delta_{n})(1-\delta_{x})$ introduced in Theorem 4, for a given $K\in\mathcal{S}$, then the estimated gradient $\hat{\nabla}C(K)$ from Algorithm 3 enjoys the following bound:
$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq 1-\delta,$ (39)
@@
$r$ $\leq r_{\max}(C(K),\epsilon_{r}),l\geq l_{\min}(C(K),\epsilon_{l}),$ (40a)
$n_{b}$ $\geq\max(N_{1},N_{3}(\hat{b}_{s}(x_{0}^{(k)}))$ (40b)
where the detailed expression of $N_{3}$ is given in (105) in Appendix C.7 and $r_{\max}$, $l_{\min}$ and $N_{1}$ were introduced in Theorem 4. Moreover, let the desired probability $\delta_{v}\in(0,1)$ be expressed as $\delta_{v}=1-(1-\tilde{\delta}_{v})(1-\tilde{\delta}_{x})$. If the number of rollouts $n_{v}$ to estimate the baseline function (as defined in Algorithm 2) satisfies
$n_{v}\geq\tilde{n}_{\min}(C(K),l,\tilde{\delta}_{v},\tilde{\delta}_{x}),$ (41)
@@
$\mathbb{P}\left\{N_{2}\geq N_{3}\right\}\geq 1-\delta_{v},$ (42)
@@
where $N_{2}$ was first introduced in (30b) in Theorem 4 and the detailed expression of $\tilde{n}_{\min}$ is provided in (110) in Appendix C.7.
@@
The proof of Theorem 6 is given in Appendix C.7. Based on Theorem 6, we can compare the performance of Algorithm 1 without variance reduction and Algorithm 3 with variance reduction. The exploration radius $r_{\max}$ and rollout length $l_{\min}$ (40a) remain the same for both algorithms. The number of rollouts (40b) is determined by the maximum of two arguments. $N_{1}$ remains unchanged across the two algorithms, as it depends on the smoothing function (see (67))and the original function. From the second part of Theorem 6 ((42)), it then follows that the required number of rollouts is at least non-increasing when the baseline function estimates are sufficiently accurate. As noted in Theorem 4, $\delta_{x}$ represents the probability that the data samples remain bounded. Increasing $\delta_{x}$ leads to a larger bound for the data samples. This upper bound is utilized in the matrix concentration inequality to determine the required number of rollouts. From (94a) and (94b), we observe that, for large $\Sigma_{w}$ and thus large noise, the required number of rollouts is determined by $N_{2}$ (in Algorithm 1) and $N_{3}$ (in Algorithm 3). In other words, the proposed variance reduction gives a provable reduction of the number of rollouts (and thus of samples) for scenarios with large noise, which is intuitive and expected.
4.3 Policy gradient descent with Adaptive Step Size
@@
Building on the gradient uncertainty quantified in the previous section, we now proceed to study the convergence of the model-free policy gradient descent algorithm.
Algorithm 4 Model-free policy gradient descent with adaptive step size
[…]
  1. Compute the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ based on Theorem 4 or 6 to achieve the accuracy stated in Theorem 7.
  2. Use Algorithm 1 or 3 to estimate the gradient $\hat{\nabla}C(\hat{K}_{i})$.
   3. Update the gradient as $\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i})$, with $\eta_{i}\leq h_{\mathrm{PGD}}(C(\hat{K}_{i}))$.
end for
Theorem 7 (Model-free policy gradient descent with Adaptive Step Size)
Suppose the initial $\hat{K}_{0}\in\mathcal{S}$, and consider the policy gradient descent with adaptive step size:
$\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i}),\quad\forall i\in\mathbb{Z}_{+},$ (43)
@@
where $\hat{\nabla}C(\hat{K}_{i})$ is the gradient estimate from Algorithm 1 and $0<\eta_{i}\leq h_{\mathrm{PGD}}(C(\hat{K}_{i}))$, where the function $h_{\mathrm{PGD}}$ was introduced in Theorem 1. Given any accuracy $\epsilon>0$, and $\sigma\in(0,1)$, define ${\eta_{\mathrm{PGD}}}:=\inf\limits_{i}\eta_{i}$ and the number of iterations $n_{\mathrm{PGD}}$:
$n_{\mathrm{PGD}}=\frac{\lVert\Sigma_{K^{*}}\rVert}{2(1-\sigma)\eta_{\mathrm{PGD}}\lambda_{1}(R)\lambda^{2}_{1}(\Sigma_{w})}\log\frac{C(\hat{K}_{0})-C(K^{*})}{\epsilon}.$
Given any probability $\delta\in(0,1)$ satisfying $\delta n_{\mathrm{PGD}}\in(0,1)$, assume the estimation error of the gradient $\hat{\nabla}C(\hat{K}_{i})$ satisfies:
@@
$\mathbb{P}\left\{\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{h_{C}(C(\hat{K}_{i}))\lVert\Sigma_{K^{*}}\rVert}\right\}\geq 1-\delta,$
@@
where $h_{C}$ was introduced in Lemma 4. Then for any $C(\hat{K}_{i})\geq C(K^{*})+\epsilon$, the following inequality holds:
$\mathbb{P}\left\{C(\hat{K}_{i+1})-C(K^{*})\leq\gamma_{i}(C(\hat{K}_{i})-C(K^{*}))\Big{|}C(\hat{K}_{i})\right\}\geq 1-\delta,$
@@
As a result, the policy gradient descent method enjoys the following performance bound:
$\mathbb{P}\left\{\min_{i\in[0,n_{\mathrm{PGD}}]}C(\hat{K}_{i})-C(K^{*})\leq\epsilon\right\}\geq 1-\delta n_{\mathrm{PGD}}.$
To achieve the desired accuracy of the estimates stated in Theorem 7, the corresponding values of $l_{i},r_{i},n_{i}$, as defined in Algorithm 4, should be selected based on Theorem 4. Theorem 7 ensures that the cost will converge to the optimal solution with a predefined accuracy gap $\epsilon$, meaning $C(K^{*})+\epsilon$ will be reached with a certain probability. However, due to the estimation error in the gradient, further improvements in cost stop when the realized cost $C(\hat{K}_{i})\leq C(K^{*})+\epsilon$. To achieve higher accuracy, it is necessary to minimize the gradient estimation error. Nevertheless, the algorithm can only converge to the optimal when the gradient estimation is exact, which is not possible with a finite number of rollouts. The interpretation of the $\sigma$ is given in Remark 3. Regarding the step size $\eta_{i}$ stated in Theorem 7, from the expression of $h_{\mathrm{PGD}}$ given in (62), we observe that as the noise level $\Sigma_{w}$ increases, the step size $\eta_{i}$ decreases, reflecting the observation analyzed in Section 3.4.
In Algorithm 4, the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ are determined adaptively to the cost $C(\hat{K}_{i})$, i.e., they are computed online inside the for-loop. This design is aligned with Theorem 7, which shows that the cost function $C(\hat{K}_{i})$ decreases probabilistically with each iteration. As the cost $C(\hat{K}_{i})$ decreases, this leads to a reduction in the required rollouts $n_{i}$ and rollout length $l_{i}$, while the exploration radius $r_{i}$ increases. Alternatively, the algorithm can be modified to compute these parameters offline, where for all iterations $i$, $n_{i},l_{i},r_{i}$ are calculated once based on the initial cost $C(\hat{K}_{0})$.
4.4 Natural Policy Gradient with Adaptive Step Size
Analogous to the model-free policy gradient descent method, we now turn to analyzing the convergence of the natural policy gradient method.